OpenStudy (anonymous):

Find a solution to the question (2/x)-(5/x^2)=-4.

OpenStudy (anonymous):

\[\frac{ 2 }{ x }-\frac{ 5 }{ x^2 }=-4\] multiply by x^2 \[2x-5=-4x^2\] \[4x^2+2x-5=0\] find x

OpenStudy (anonymous):

You are a magical being.

OpenStudy (anonymous):

Do you have any idea what a graph for that would look like